# The Voronoi Spherical CDF for Lattices and Linear Codes

### v2

@@ Appendix D Normalized $p$th Moment
${\mathbb{E}}\|U_{L}\|^{p}_{p}=\int_{0}^{\infty}\Pr(\|U_{L}\|^{p}_{p}>r)dr=\int_{0}^{\infty}\Pr(\|U_{L}\|_{p}>r^{1/p})dr.$ (161)
@@
${\mathbb{E}}[Y]$ $\leq\frac{n+p}{n}\frac{1}{{\mathrm{sinc}}(p/n)}-1\leq\left(1+\frac{p}{n}\right)\left(1+2\left(\frac{p}{n}\right)^{2}\right)-1$
@@ Appendix E Proofs of claims for linear codes
${r_{\mathrm{pack}}}({\mathcal{C}})=\max\{r\penalty 10000\ :\penalty 10000\ {\mathcal{B}}_{n,r}\subset\mathcal{V}_{\mathcal{C}}\}=\max\{r\penalty 10000\ :\penalty 10000\ Q_{{\mathcal{C}}}(r)=2^{k-n}V_{n,r}\},$ (177)
@@
${\mathbb{E}}|U_{L}|=\sum_{r=0}^{n}\Pr(|U_{\mathcal{C}}|>r)=\sum_{r=0}^{n}1-\Pr(|U_{\mathcal{C}}|\leq r)=\sum_{r=0}^{n}1-Q_{\mathcal{C}}(r).$ (179)
@@
$=\left(\frac{p}{1-p}+\frac{1-2p}{1-p}\right)\varphi_{p}(n)\left|{\mathcal{K}}\cap{\mathcal{B}}_{n,n}\right|+\frac{1-2p}{1-p}\sum_{r=0}^{n-1}\varphi_{p}(r)\left|{\mathcal{K}}\cap{\mathcal{B}}_{n,r}\right|$ (184)
@@
$=|{\mathcal{K}}|\cdot\frac{p}{1-p}\cdot p^{n}+|{\mathcal{K}}|\cdot\frac{1-2p}{1-p}\sum_{r=0}^{n}{{n}\choose{r}}p^{r}(1-p)^{n-r}\frac{\left|{\mathcal{K}}\cap{\mathcal{B}}_{n,r}\right|/|{\mathcal{K}}|}{{{n}\choose{r}}}$ (186)
@@
$\gamma_{r}$ $=\frac{{{n}\choose{r+1}}}{V_{n,r}}=\frac{{{n}\choose{r+1}}}{{{n}\choose{r}}+V_{n,r-1}}=\frac{{{n}\choose{r+1}}}{{{n}\choose{r}}\left(1+\frac{1}{\gamma_{r-1}}\right)}$ (191)
@@
$\frac{x_{r+1}}{x_{r}}=\frac{V_{n,r+1}}{V_{n,r}}=1+\frac{{{n}\choose{r+1}}}{V_{n,r}}=1+\gamma_{r}\geq\beta_{m},\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall 0<r\leq m.$ (193)
@@
$S_{1}$ $\leq\sum_{r=0}^{{r_{\mathrm{eff}}}(n,k)-1}x_{r}\leq\sum_{r=0}^{{r_{\mathrm{eff}}}(n,k)-1}\beta_{m}^{-({r_{\mathrm{eff}}}(n,k)-r)}<\sum_{j=0}^{\infty}\beta_{m}^{-j}=\frac{1}{1-\beta_{m}^{-1}}$ (200)
$S_{2a}$ $\leq\sum_{r={r_{\mathrm{eff}}}(n,k)}^{m}x^{-1}_{r}\leq\sum_{r={r_{\mathrm{eff}}}(n,k)}^{m}\beta_{m}^{-(r-{r_{\mathrm{eff}}}(n,k))}<\sum_{j=1}^{\infty}\beta_{m}^{-j}\leq\frac{1}{1-\beta_{m}^{-1}}.$ (201)
@@
Take $m=\left\lceil\frac{n}{2}\left(h_{2}^{-1}(1-\alpha)+\frac{1}{2}\right)\right\rceil$,cwhere $h_{2}(p)=-p\log_{2}(p)-(1-p)\log_{2}(1-p)$, and $h_{2}^{-1}(\cdot)$ is its inverse restricted to $[0,1/2)$. We may assume without loss of generality that $n$ is large enough such that $\kappa{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\frac{m+1}{n}<1/2$, so that $\beta_{m}\geq\frac{1-\kappa}{\kappa}>1$, and $m-{r_{\mathrm{eff}}}(n,k)>\left\lceil\frac{\log n}{\log\beta_{m}}\right\rceil$. Substituting this into (202), we obtain

### v3

@@ Appendix D Normalized $p$th Moment
${\mathbb{E}}\|U_{L}\|^{p}_{p}$ $=\int_{0}^{\infty}\Pr(\|U_{L}\|^{p}_{p}>r)dr$ (185)
$=\int_{0}^{\infty}\Pr(\|U_{L}\|_{p}>r^{1/p})dr.$ (186)
@@
${\mathbb{E}}[Y]$ $\leq\frac{n+p}{n}\frac{1}{{\mathrm{sinc}}(p/n)}-1$ (201)
$\leq\left(1+\frac{p}{n}\right)\left(1+2\left(\frac{p}{n}\right)^{2}\right)-1$ (202)
@@ Appendix E Proofs of claims for linear codes
${r_{\mathrm{pack}}}({\mathcal{C}})$ $=\max\{r\penalty 10000\ :\penalty 10000\ {\mathcal{B}}_{n,r}\subset\mathcal{V}_{\mathcal{C}}\}$ (204)
$=\max\{r\penalty 10000\ :\penalty 10000\ Q_{{\mathcal{C}}}(r)=2^{k-n}V_{n,r}\},$ (205)
@@
${\mathbb{E}}|U_{L}|$ $=\sum_{r=0}^{n}\Pr(|U_{\mathcal{C}}|>r)$ (207)
$=\sum_{r=0}^{n}1-\Pr(|U_{\mathcal{C}}|\leq r)$ (208)
$=\sum_{r=0}^{n}1-Q_{\mathcal{C}}(r).$ (209)
@@
$=\left(\frac{p}{1-p}+\frac{1-2p}{1-p}\right)\varphi_{p}(n)\left|{\mathcal{K}}\cap{\mathcal{B}}_{n,n}\right|$
$+\frac{1-2p}{1-p}\sum_{r=0}^{n-1}\varphi_{p}(r)\left|{\mathcal{K}}\cap{\mathcal{B}}_{n,r}\right|$ (214)
@@
$=|{\mathcal{K}}|\cdot\frac{p}{1-p}\cdot p^{n}$
$+|{\mathcal{K}}|\cdot\frac{1-2p}{1-p}\sum_{r=0}^{n}{{n}\choose{r}}p^{r}(1-p)^{n-r}\frac{\left|{\mathcal{K}}\cap{\mathcal{B}}_{n,r}\right|/|{\mathcal{K}}|}{{{n}\choose{r}}}$ (216)
@@
$\gamma_{r}$ $=\frac{{{n}\choose{r+1}}}{V_{n,r}}=\frac{{{n}\choose{r+1}}}{{{n}\choose{r}}+V_{n,r-1}}=\frac{{{n}\choose{r+1}}}{{{n}\choose{r}}\left(1+\frac{1}{\gamma_{r-1}}\right)}$
@@
$\frac{x_{r+1}}{x_{r}}$ $=\frac{V_{n,r+1}}{V_{n,r}}$ (222)
$=1+\frac{{{n}\choose{r+1}}}{V_{n,r}}$ (223)
$=1+\gamma_{r}$ (224)
$\geq\beta_{m},\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall 0<r\leq m.$ (225)
@@
$S_{1}$ $\leq\sum_{r=0}^{{r_{\mathrm{eff}}}(n,k)-1}x_{r}\leq\sum_{r=0}^{{r_{\mathrm{eff}}}(n,k)-1}\beta_{m}^{-({r_{\mathrm{eff}}}(n,k)-r)}<\sum_{j=0}^{\infty}\beta_{m}^{-j}$
$=\frac{1}{1-\beta_{m}^{-1}}$ (232)
$S_{2a}$ $\leq\sum_{r={r_{\mathrm{eff}}}(n,k)}^{m}x^{-1}_{r}\leq\sum_{r={r_{\mathrm{eff}}}(n,k)}^{m}\beta_{m}^{-(r-{r_{\mathrm{eff}}}(n,k))}<\sum_{j=1}^{\infty}\beta_{m}^{-j}$
$\leq\frac{1}{1-\beta_{m}^{-1}}.$ (233)
@@
Take $m=\left\lceil\frac{n}{2}\left(h_{2}^{-1}(1-\alpha)+\frac{1}{2}\right)\right\rceil$, where $h_{2}(p)=-p\log_{2}(p)-(1-p)\log_{2}(1-p)$, and $h_{2}^{-1}(\cdot)$ is its inverse restricted to $[0,1/2)$. We may assume without loss of generality that $n$ is large enough such that $\kappa{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\frac{m+1}{n}<1/2$, so that $\beta_{m}\geq\frac{1-\kappa}{\kappa}>1$, and $m-{r_{\mathrm{eff}}}(n,k)>\left\lceil\frac{\log n}{\log\beta_{m}}\right\rceil$. Substituting this into (234), we obtain